How many possible samples n=2 can be drawn from a population of size 14?

1
Expert's answer
2022-05-12T03:09:31-0400

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{14}C_2=91."

The number of possible samples which can be drawn with replacement is "N^n=14^2=196."